Andrew has a box which contains 4 pink blocks, 5 yellow blocks and 6 green blocks. He picks one block, records its color and puts it back in the box. He then draws another block. What is the probability of taking out a yellow block followed by the pink block?
Answer #1

Solution:-

=> P(Yellow) = 5/(4+5+6) = 5/16

=> P(Pink) = 4/16 = 1/4


=> Probability = (5/16)*(1/4) = 5/64 = 0.0781